OpenStudy (experimentx):

Prove:- \[ \lim_{N\rightarrow\infty}\sum_{k=1}^N\left(\frac{k-1}{N}\right)^N = {1 \over e - 1}\]

OpenStudy (zzr0ck3r):

I only know how to prove limits of sequences not series:(

OpenStudy (zzr0ck3r):

what a cool series though..

OpenStudy (experimentx):

yep!! this is very interesting!!

OpenStudy (zzr0ck3r):

god we really dont know what we have on our hands with this math stuff.

OpenStudy (anonymous):

I messed up on my first one... \[\lim_{N\rightarrow\infty}\left(\frac{k-1}{N}\right)^N=\lim_{N\rightarrow\infty}\left(1+\frac{k-1-N}{N}\right)^N=e^{k-1-N}.\]

OpenStudy (experimentx):

you missed the sum. the approach is +ve ... BRB in 3 hrs

OpenStudy (anonymous):

And you can use a geometric series from there to get \(\displaystyle \frac{e^{-1}}{1-e^{-1}}\)

OpenStudy (anonymous):

I know, it's not very rigorous. I've only learned as much analysis as I've taught myself, so I'm still not very good with proofs...

OpenStudy (experimentx):

change of limits on the sum is the trick!!

OpenStudy (experimentx):

sorry.. what do we call that ... k ... counter as in loop

OpenStudy (anonymous):

Are you allowed to take the limit through the sum, as I have done?

OpenStudy (anonymous):

For example, to go from:\[\large\lim_{x\rightarrow\infty}\sum_{k=0}^\infty f(x)\overset{?}{=}\sum_{k=0}^\infty \lim_{x\rightarrow\infty}f(x)\]

OpenStudy (experimentx):

this is equivalent to lim n->inf (1 - 1/n)^n + (1 - 2/n)^n + (1 - 3/n)^n + ....

OpenStudy (experimentx):

let's put a new counter m = N - k + 1 <-- this counts from opposite.

OpenStudy (anonymous):

Yeah, that's what I did too. I didn't know it was called a counter, though.

OpenStudy (experimentx):

i don't know what it is called either ... may be some incremental variable.

OpenStudy (experimentx):

in programming ... in loop ... it's called counter... :D

OpenStudy (anonymous):

lol

OpenStudy (experimentx):

gotta go ... bye